# Rare Event Detection in Imbalanced Multi-Class Datasets Using an Optimal MIP-Based Ensemble Weighting Approach

### v1

@@ Introduction
Rare event detection in cyber-physical systems (CPS) is vital for maintaining the safety and reliability of critical infrastructures, such as water distribution networks, power grids, and transportation networks (Taheri et al. 2024; Tertytchny, Nicolaou, and Michael 2020). Critical CPS often generate imbalanced multi-class datasets where some classes have significantly fewer instances than others, due to the infrequent occurrence of abnormal events (e.g., faults or attacks) compared to normal activities (He and Garcia 2009; Yin et al. 2020). Class imbalance poses several challenges, such as biased model performance towards the majority classes, poor generalization to minority classes, and thus increased difficulty in detecting rare events (Shyalika, Wickramarachchi, and Sheth 2023). These issues are further amplified by the critical nature of CPS, where misclassifying low-probability but high-impact events may lead to safety risks or system failures. Consequently, addressing these challenges is crucial for effective rare event detection.
@@
Ensemble learning approaches, especially weighted voting techniques, are commonly employed to handle imbalanced multi-class datasets (Khan, Chaudhari, and Chandra 2024). The aim of such methods is to improve overall performance and robustness, by leveraging the strengths of diverse classification algorithms. Specifically, a weighted voting ensemble model combines the predictions of multiple base classifiers, using different weights based on their predictive performance. Higher weights are assigned to better-performing classifiers, either at an overall classifier level or on a per class basis (Dogan and Birant 2019). However, existing weighting schemes often fail to optimally balance the contribution of each classifier, especially in the presence of a highly skewed class distribution. Moreover, since smaller ensembles are typically easier to manage and less computationally demanding, and given that minimizing latency necessitates performing rare event detection at the edge of the network (often on devices with limited resources), adaptable approaches are required to seamlessly and optimally select a predefined number of classifiers from a larger set while assigning their weights (Danso et al. 2022).
@@
Aiming to fill this gap, we propose an optimal, adaptable, and efficient mixed integer programming (MIP) weighting scheme for weighted voting ensembles, suitable for imbalanced multi-class datasets. Our approach optimally assigns weights to classifiers per class based on their performance, leveraging their individual capabilities while selecting the predefined number of classifiers from a given set. To prevent overfitting and enhance generalization in weight assignment, we incorporate into our method the elastic net regularization technique (Zou and Hastie 2005), applying it beyond its traditional use in the training and validation phase. The proposed weighting scheme not only improves performance across all classes, but it also ensures computational efficiency, addressing the limitations of existing approaches and thus providing a more robust solution for rare event detection in critical CPS.
@@
To our knowledge, in contrast to our approach, none of the existing ensemble weighting schemes is intrinsically adaptable to varied ensemble sizes and capable of optimally assigning classifier weights in imbalanced multi-class datasets, while attaining generalization and efficiency.
@@
Weighted voting ensemble methods typically apply weights per classifier or per classifier and class, based on classifier performance in the training and validation phase (Maheshwari et al. 2022). Among the fundamental techniques, UW-PC (uniform weights per classifier) and UW-PCC (uniform weights per classifier and class) assign equal weights to all classifiers and classifier-class pairs, respectively, and commonly serve as baselines in comparative studies (Dong et al. 2020). On the other hand, WA-PC (weighted average based on normalized accuracy per classifier) and WA-PCC (weighted average based on normalized accuracy per classifier and class) apply weights based on the normalized validation accuracy of each classifier or classifier-class pair, respectively (Sagi and Rokach 2018). Both techniques assign greater weights to classifiers or classifier-class pairs that exhibit better performance. DE (differential evolution) is an evolutionary algorithm that determines classifier weights in an iterative refining manner (Storn and Price 1997; Zhang et al. 2014, 2024). In contrast, BMA (Bayesian model averaging) assigns weights to classifier-class pairs based on their Bayesian inference posterior probabilities, which reflect the likelihood of each classifier being accurate given the observed data (Raftery, Madigan, and Hoeting 1997; Wang et al. 2022).
However, these weighting schemes do not optimally assign classifier weights, are not easily adaptable to a varied number of classifiers, and fail to address highly uneven class distributions while achieving generalization. In contrast, our MIP approach attains generalization in weight calculation through elastic net regularization, which is a regularized regression technique. It linearly combines the widely used L1 and L2 penalties of the lasso and ridge methods, respectively, to leverage the advantages of both strategies (Zou and Hastie 2005). This dual technique can lead to a weight assignment that is both sparse (simpler and less likely to overfit) and robust (less sensitive to training data noise) (Giesen et al. 2019).
@@
    Test & Evaluation: The calculated weights in $\mathbf{W}$ are used to evaluate the performance of the resulting ensemble model comprising the selected $K$ classifiers.
@@
Considering the sets $\mathcal{C}$ and $\mathcal{E}$, matrices $\mathbf{V}$ and $\mathbf{W}$, and parameter $K$, we formulate the problem as a MIP model.
@@
We utilize the following variables:
@@
The three phases of our methodology were implemented using a server running CentOS 7.9, equipped with an Intel Xeon Gold 6240 processor @ 2.6 GHz, 400 GB of RAM, and two NVIDIA Tesla V100 GPUs, each with 32 GB of dedicated memory. The software environment included Python 3.12 with scipy 1.14.0, scikit-learn (sklearn) 1.5.1, NumPy 2.0, pandas 2.2.2, and Gurobi Optimizer 11.0.3 with its Python interface gurobipy. We utilized Gurobi Optimizer in phase 2 to solve our MIP-formulated model, as it is widely used to optimally solve similar problems (Gurobi Optimization, LLC 2024). We split each examined dataset into 80% for training and validation (60% for training and 20% for validation), used in phase 1, and 20% for testing, used in phase 3. For the stratified $k$-fold cross-validation in phase 1 we employed $k=5$ folds (McTavish et al. 2022). To fine-tune parameters $\lambda$ and $\alpha$ in (1), we conducted preliminary experimentation, which revealed that a narrow range of values ($0.95\leq\lambda\leq 1$ and $0.80\leq\alpha\leq 0.87$) consistently yielded the best performance across different datasets. The process involved incrementing/decrementing each parameter and terminating once the performance started to deteriorate. Consequently, to fine-tune $\lambda$ and $\alpha$ for each dataset, we explored only these particular ranges. Our fine-tuning resulted in the following values for $(\lambda,\alpha)$ for each of the four examined datasets: $(0.95,0.85)$, $(0.96,0.80)$, $(1.00,0.82)$, and $(0.95,0.86)$, respectively.
@@
We experimented with four representative datasets for rare event detection in critical CPS, featuring multiple classes with varied imbalance: [D1] LeakDB (Vrachimis et al. 2018), [D2] NSL-KDD (University of New Brunswick, Canada 2024b), [D3] SG-MITM (Elrawy et al. 2023), and [D4] CIC-IDS2017 (University of New Brunswick, Canada 2024a). Benchmarks D2 and D4 pertain to intrusion detection and cybersecurity threats in simulated network environments, resembling real-world network traffic captures. They are widely used to investigate anomaly detection in CPS (Zhou et al. 2020; Danso et al. 2022). On the other hand, D1 involves leakages (faults) in a simulated real-world water distribution network under varying conditions, while D3 focuses on man-in-the-middle (MITM) attacks targeting client–server protocols in the private area network (PAN) of a smart grid CPS, represented by an experimental testbed. Table 1 shows the description, number of instances, and the percentage distribution of each normal/abnormal class in datasets D1–D4. Furthermore, it showcases the imbalance ratio $\rho$ for each dataset, which is defined as the ratio of the number of instances $\mu$ in the majority class to the number of instances $\nu$ in the minority class (Wang et al. 2021). Considering that a dataset is highly imbalanced when $\mu\gg\nu$, D1 and D2 are highly imbalanced, whereas D3 and D4 are moderately imbalanced (Liu et al. 2020).
@@
Figures 25 demonstrate the percentage improvement achieved in balanced accuracy (primary metric) by the proposed MIP weighting scheme over all existing approaches for datasets D1–D4, respectively, for $K=2$ to $K=8$ classifiers selected from set $\mathcal{C}$. In addition, Tables 25 showcase the percentage improvement attained with respect to the secondary metrics (macro-averaged precision, recall, and F1-score) in each case. The average improvement per dataset is also shown in Table 6. Notably, MIP consistently outperformed the other techniques across all evaluation metrics, ensemble sizes, and datasets, demonstrating superior generalization and robustness. Specifically, MIP yielded an increase in balanced accuracy ranging from 0.99% to 7.31%, with an overall average of 4.53% across all datasets and ensemble sizes. Moreover, it attained an overall average increase of 4.63%, 4.60%, and 4.61% in macro-averaged precision, recall, and F1-score, respectively.
Furthermore, it can be observed that the improvement provided in all performance metrics by our approach over the other schemes generally increased as the ensemble size $K$ increased. This reveals that MIP benefits more from a larger number of classifiers compared to the other methods. By optimally assigning weights to classifier-class pairs, our approach enhances the ability of the ensemble to leverage the diverse capabilities of each classifier, leading to a better performance as the number of classifiers grows. On the other hand, the performance improvement was lower for the highly imbalanced datasets D1 and D2 (which had a higher imbalance ratio), compared to the moderately imbalanced datasets D3 and D4 (which had a lower imbalance ratio). This highlights the negative impact of class imbalance on the ensemble performance under all weighting schemes. However, as demonstrated in Table 6, the performance gains for the highly imbalanced datasets were still significant and comparable to those observed for the moderately imbalanced datasets. As reported by the Gurobi solver, the number of variables and constraints required for each dataset by our MIP approach are shown in Table 6.

### v2

@@ Introduction
Rare event detection in cyber-physical systems (CPS) is vital for maintaining the safety and reliability of critical infrastructures, such as water distribution networks and power grids (Taheri et al. 2024; Tertytchny, Nicolaou, and Michael 2020). Critical CPS often generate imbalanced multi-class datasets where some classes have significantly fewer instances than others, due to the infrequent occurrence of abnormal events (e.g., faults or attacks) compared to normal activities (Yin et al. 2020). Class imbalance poses several challenges, such as biased model performance towards the majority classes, poor generalization to minority classes, and thus increased difficulty in detecting rare events (Shyalika, Wickramarachchi, and Sheth 2023). These issues are further amplified by the critical nature of CPS, where misclassifying low-probability but high-impact events may lead to safety risks or system failures. Consequently, addressing these challenges is crucial for effective rare event detection.
@@
Ensemble learning approaches, especially weighted voting techniques, are commonly employed to handle imbalanced multi-class datasets (Khan, Chaudhari, and Chandra 2024). The aim of such methods is to improve overall performance and robustness, by leveraging the strengths of diverse classification algorithms. Specifically, a weighted voting ensemble model combines the predictions of multiple classifiers, using different weights based on their predictive performance. Higher weights are assigned to better-performing classifiers, either at an overall classifier level or on a per class basis (Dogan and Birant 2019). However, existing weighting schemes often fail to optimally balance the contribution of each classifier, especially in the presence of a highly skewed class distribution. Moreover, since smaller ensembles are typically less computationally demanding, and given that rare event detection in CPS is often performed at the network edge on resource-constrained devices, adaptable approaches are required to seamlessly and optimally select a predefined number of classifiers from a larger set while assigning their weights (Danso et al. 2022).
@@
Aiming to fill this gap, we propose an optimal, adaptable, and efficient mixed integer programming (MIP) weighting scheme for weighted voting ensembles, suitable for rare event detection in imbalanced multi-class datasets. Our approach optimally assigns weights to classifiers per class based on their performance, leveraging their individual capabilities while selecting the predefined number of classifiers from a given set. To prevent overfitting and enhance generalization in weight assignment, we incorporate into our method the elastic net regularization technique (Zou and Hastie 2005), applying it beyond its traditional use in the training and validation phase. The proposed weighting scheme aims to enhance performance across all classes while ensuring computational efficiency, effectively addressing the limitations of existing approaches. Our method optimizes weights based on empirical accuracy results, avoiding simplifying assumptions typically required by theoretical analysis, such as independent and identically distributed instances, which may not hold in realistic scenarios involving class imbalance and diverse classifier behavior (Opitz and Maclin 1999).
@@
To our knowledge, in contrast to our approach, none of the existing ensemble weighting schemes are intrinsically adaptable to varied ensemble sizes and capable of optimally assigning classifier weights for imbalanced multi-class datasets, while attaining generalization and efficiency.
@@
Weighted voting ensemble methods typically apply weights per classifier or per classifier and class, based on classifier performance in the training and validation phase (Maheshwari et al. 2022). Among the fundamental techniques, UW-PC (uniform weights per classifier) and UW-PCC (uniform weights per classifier and class) assign equal weights to all classifiers and classifier-class pairs, respectively, and commonly serve as baselines in comparative studies (Dong et al. 2020). On the other hand, WA-PC (weighted average based on normalized accuracy per classifier) and WA-PCC (weighted average based on normalized accuracy per classifier and class) apply weights based on the normalized accuracy of each classifier or classifier-class pair, respectively (Sagi and Rokach 2018). Both techniques assign greater weights to classifiers or classifier-class pairs that exhibit better performance. DE (differential evolution) is an evolutionary algorithm that determines classifier weights in an iterative refining manner (Storn and Price 1997; Zhang et al. 2014, 2024). In contrast, BMA (Bayesian model averaging) assigns weights to classifier-class pairs based on their Bayesian inference posterior probabilities, which reflect the likelihood that each classifier is accurate given the observed data (Raftery, Madigan, and Hoeting 1997; Wang et al. 2022).
However, these weighting schemes do not optimally assign classifier weights, are not easily adaptable to a varied number of classifiers, and fail to address highly uneven class distributions while achieving generalization. In contrast, our MIP approach attains generalization in weight calculation through elastic net regularization, which is a regularized regression technique. It linearly combines the widely used L1 and L2 penalties of the lasso and ridge methods, respectively, to leverage the advantages of both strategies (Zou and Hastie 2005). This dual technique can lead to a weight assignment that is both sparse (less likely to overfit) and robust (less sensitive to training data noise) (Giesen et al. 2019).
@@
    Test & Evaluation: The calculated weights in $\mathbf{W}$ are assigned to the selected $K$ classifiers to evaluate the performance of the resulting ensemble model.
The main problem addressed—determining the optimal class-based weight assignment for classifiers in a weighted voting ensemble model—corresponds to phase 2.
@@
To solve the examined problem in phase 2, we formulate it as a MIP model by defining its decision variables, objective function, and constraints.
@@
The following variables are used to select $K$ classifiers and to determine their optimal weights:
@@
The time complexity of MIP models depends on both the problem size and the solver used. MIP solvers typically employ proprietary algorithms with undisclosed implementation details, and thus their time complexity cannot be easily derived. However, the computational cost of the proposed MIP approach can be considered proportional to the problem size $n\,m$, where $n$ (number of classifiers) is small in practice (Dogan and Birant 2019). Consequently, the computational cost grows approximately linearly with respect to $m$ (number of classes). In CPS rare event detection, $m$ is typically small (Vrachimis et al. 2018), as the aim is to distinguish between normal and rare abnormal states rather than to conduct a fine-grained classification. It is noted that our MIP model is independent of the dataset size. These factors ensure that the runtime of our MIP method remains reasonable, as demonstrated in our experimental results.
@@
The three phases of our methodology were implemented using a server running CentOS 7.9, equipped with an Intel Xeon Gold 6240 processor @ 2.6 GHz and 400 GB of RAM. The software environment included Python 3.12 with scipy 1.14.0, scikit-learn 1.5.1, NumPy 2.0, pandas 2.2.2, and Gurobi Optimizer 11.0.3 with its Python interface gurobipy. We utilized the Gurobi solver in phase 2 to solve our MIP-formulated model, as it is widely used to optimally solve similar optimization problems (Gurobi, LLC 2024). We split each examined dataset into 80% for training and validation (60% for training and 20% for validation), used in phase 1, and 20% for the test set, used in phase 3. For the stratified $k$-fold cross-validation in phase 1 we employed $k=5$ folds (McTavish et al. 2022). To fine-tune parameters $\lambda$ and $\alpha$ in (1), we incremented/decremented each parameter, terminating adjustments when performance began to deteriorate. This fine-tuning resulted in the following values for $(\lambda,\alpha)$ for each of the four examined datasets: $(0.95,0.85)$, $(0.96,0.80)$, $(1.00,0.82)$, and $(0.95,0.86)$, respectively.
@@
We experimented with four representative datasets for rare event detection in critical CPS, featuring multiple classes with varied imbalance: [D1] LeakDB (Vrachimis et al. 2018), [D2] NSL-KDD (Tavallaee et al. 2009), [D3] SG-MITM (Elrawy et al. 2023), and [D4] CIC-IDS2017 (Sharafaldin, Lashkari, and Ghorbani 2018). Benchmarks D2 and D4 pertain to intrusion detection in simulated network environments, resembling real-world network traffic captures. They are widely used to investigate anomaly detection in CPS (Zhou et al. 2020; Danso et al. 2022). On the other hand, D1 involves leakages (faults) in a simulated real-world water distribution network under varying conditions, while D3 focuses on man-in-the-middle (MITM) attacks targeting client-server protocols in the private area network (PAN) of a smart grid CPS, represented by an experimental testbed. Table 1 shows the description, number of instances, and the percentage distribution of each normal/abnormal class in datasets D1–D4. Furthermore, it showcases the imbalance ratio $\rho$ for each dataset, which is defined as the ratio of the number of instances $\mu$ in the majority class to the number of instances $\nu$ in the minority class (Wang et al. 2021). Considering that a dataset is highly imbalanced when $\mu\gg\nu$, D1 and D2 are highly imbalanced, whereas D3 and D4 are moderately imbalanced (Liu et al. 2020).
@@
Figures 25 demonstrate the percentage improvement achieved in balanced accuracy (primary metric) by the proposed MIP weighting scheme over all existing approaches for datasets D1–D4, respectively, for $K=2$ to $K=8$ classifiers selected from set $\mathcal{C}$. Tables 25 showcase the percentage improvement attained with respect to the secondary metrics (macro-averaged precision, recall, and F1-score) in each case. The average improvement per dataset is also shown in Table 6. Notably, MIP consistently outperformed the other techniques across all evaluation metrics, ensemble sizes, and datasets, demonstrating superior generalization and robustness. Specifically, MIP yielded an increase in balanced accuracy ranging from 0.99% to 7.31%, with an overall average of 4.53% across all datasets and ensemble sizes. Moreover, it attained an overall average increase of 4.63%, 4.60%, and 4.61% in macro-averaged precision, recall, and F1-score, respectively. In addition to these metrics, MIP also outperformed all other schemes with respect to the macro-averaged area under the precision-recall curve (AUPRC), as detailed in the provided supplementary material.
Furthermore, it can be observed that the improvement provided in all performance metrics by our approach over the other schemes generally increased as the ensemble size $K$ increased. This reveals that MIP benefits more from a larger number of classifiers compared to the other methods. By optimally assigning weights to classifier-class pairs, our approach enhances the ability of the ensemble to leverage the diverse capabilities of each classifier, leading to a better performance as the number of classifiers grows. On the other hand, the performance improvement was lower for the highly imbalanced datasets D1 and D2 (which had a higher imbalance ratio), compared to the moderately imbalanced datasets D3 and D4 (which had a lower imbalance ratio). This highlights the negative impact of class imbalance on the ensemble performance under all weighting schemes. However, as demonstrated in Table 6, the performance gains for the highly imbalanced datasets were still significant and comparable to those observed for the moderately imbalanced datasets. The effectiveness and robustness of our method under different cases of class imbalance are further demonstrated through the ablation studies included in the supplementary material. As reported by the Gurobi solver, the number of variables and constraints required for each dataset by our MIP approach are shown in Table 6.